OpenStudy (anonymous):

Find a cubic function with the given zeros. -7, 5, -3

OpenStudy (anonymous):

you know these are zeros this means (x-a)(x-b)(x-c)=0 x= -7,5,-3 solve for a,b,c

OpenStudy (anonymous):

i did that but i wasnt getting the answers that were shown

OpenStudy (anonymous):

can you show me what you got?

OpenStudy (anonymous):

x^3 +2x^2-35x

OpenStudy (amistre64):

35*3 is not 35 ...

OpenStudy (anonymous):

you're missing a constant can you write it in the form of (x-a)(x-b)(x-c)=0?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

these are my choices f(x) = x3 + 5x2 - 29x - 105 f(x) = x3 + 5x2 - 29x + 105 f(x) = x3 + 5x2 + 29x - 105 f(x) = x3 - 5x2 - 29x - 105

OpenStudy (anonymous):

i would ignore those for now (x+7)(x-5)(x+3)=0 lets do this step by step (x+7)(x-5)=x(x-5)+7(x-5)=???

OpenStudy (anonymous):

ok...?

OpenStudy (anonymous):

solve this (x+7)(x-5)=x(x-5)+7(x-5)=

OpenStudy (anonymous):

x2 - 5x + 7x + 21 x2 + 2x +21

OpenStudy (anonymous):

7(x-5)= 7x-35, not 7x+21

OpenStudy (anonymous):

jk i was doing 7(x+3)

OpenStudy (anonymous):

... (x+7)(x-5)= x^2+2x-35 now we multiply x+3 to this

OpenStudy (anonymous):

(x^2+2x-35)(x+3)=??? (x^2+2x-35)(x+3)=x(x^2+2x-35)+3(x^2+2x-35)=???

OpenStudy (anonymous):

ah i got it!!

OpenStudy (anonymous):

thank you!!!